# Task Aware Dreamer for Task Generalization in Reinforcement Learning

### v4

@@ A.3 The Proof of Theorem 2
For any $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R}_{\mathcal{M}},\gamma)\sim\mathcal{T}$, we can use the policy $\pi$ to interact with $\mathcal{M}$ and get the trajectory $\tau=(s_{0}^{\mathcal{M}},a_{0}^{\mathcal{M}},r_{1}^{\mathcal{M}},s_{1}^{\mathcal{M}},a_{1}^{\mathcal{M}},r_{2}^{\mathcal{M}},...)$. Since the dynamic transition $\mathcal{P}$ is the same for all $\mathcal{M}$ and the policy $\pi\in\Pi_{2}$ only depends on historical states and actions, we naturally have that the distribution of all states and actions $(s_{0}^{\mathcal{M}},a_{0}^{\mathcal{M}},s_{1}^{\mathcal{M}},a_{1}^{\mathcal{M}},...)$ are the same for all $\mathcal{M}\sim\mathcal{T}$ as well as $\bar{\mathcal{M}}$. Consequently, we have
$\begin{split}&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]\\ =&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{\tau\sim\mathcal{P},\pi}[R_{\mathcal{M}}(\tau)]=\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}^{\mathcal{M}}\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[r_{t}^{\mathcal{M}}]\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\mathcal{R}_{\mathcal{M}}(s_{t}^{\mathcal{M}},a_{t}^{\mathcal{M}})]\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}[\bar{\mathcal{R}}(s_{t}^{\mathcal{M}},a_{t}^{\mathcal{M}})]\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}[R_{\bar{\mathcal{M}}}(\tau)]=J_{\bar{\mathcal{M}}}(\pi).\end{split}$ (13)
@@
$\begin{split}J_{\mathcal{T}}^{2}&=\max_{\pi\in\Pi_{2}}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=\max_{\pi\in\Pi_{2}}J_{\bar{\mathcal{M}}}(\pi)\\ &=\max_{\pi\in\Pi_{1}}J_{\bar{\mathcal{M}}}(\pi)=\max_{\pi\in\Pi_{1}}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=J_{\mathcal{T}}^{1}\\ &\leq\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[\max_{\pi\in\Pi_{1}}J_{\mathcal{M}}(\pi)\right]=J_{\mathcal{T}}^{*}.\end{split}$ (14)
Thus we have proven this result. ∎
@@ A.4 The Proof of Theorem 3
Proof.
Our proof follows some previous work (Kakade and Langford 2002; Ying et al. 2022). First, we consider the bellman equation of value function of $\pi,\pi_{\mathcal{M}}^{*}\in\Pi_{1}$ in $\mathcal{M}$ as
@@
$\begin{split}&V_{\mathcal{M},\pi}(s)-V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)\\ =&\gamma\sum_{a}\Delta\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s^{\prime})\\ +&\gamma\sum_{a}\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)\Delta V(s^{\prime})+\sum_{a}\Delta\pi(a|s)\mathcal{R}(s,a)\\ =&\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ +&\gamma\sum_{a}\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)\Delta V(s^{\prime}),\end{split}$ (15)
@@
$\begin{split}&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\Delta V(s)\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)[V_{\mathcal{M},\pi}(s)-V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)]\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ +&\gamma\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)\Delta V(s^{\prime})\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ +&\sum_{s^{\prime}}\Delta V(s^{\prime})\left[\gamma\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\pi(a|s)\mathcal{P}(s^{\prime}|s,a)\right].\end{split}$ (16)
@@
$\begin{split}\sum_{s}d_{\mathcal{M}}^{\pi}(s)&\Delta V(s)=\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ +&\sum_{s^{\prime}}\Delta V(s^{\prime})\left[d_{\mathcal{M}}^{\pi}(s^{\prime})-(1-\gamma)\mathcal{P}(s_{0}=s^{\prime})\right].\end{split}$ (17)
@@
$\begin{split}&(1-\gamma)\sum_{s^{\prime}}\Delta V(s^{\prime})\mathcal{P}(s_{0}=s^{\prime})\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a),\end{split}$ (18)
@@
[…]
@@
$\begin{split}&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi_{\mathcal{M}}^{*})-J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}(\cdot)}\left[\max_{a}Q_{\mathcal{M}}^{*}(s,a)\right.\\ &\left.-\int_{a}\pi(a|s)Q_{\mathcal{M}}^{*}(s,a)da\right]\\ \geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}(\cdot)}\left[\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\max_{a}Q_{\mathcal{M}}^{*}(s,a)\right.\\ &\left.-\max_{a}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}Q_{\mathcal{M}}^{*}(s,a)\right]\\ =&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}}\left[D_{\text{TDR}}(\mathcal{T},s)\right].\end{split}$ (20)
@@
$\begin{split}&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\max_{\pi\in\Pi_{1}}J_{\mathcal{M}}(\pi)-\max_{\pi\in\Pi_{1}}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}J_{\mathcal{M}}(\pi)\\ \geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi^{*}}}\left[D_{\text{TDR}}(\mathcal{T},s)\right],\end{split}$ (21)
Thus we have proven this result. ∎
@@ A.5 Expressive Ability of $\Pi_{3}$
$\begin{split}&\mathcal{P}(s_{t}^{l},a_{j})=\mathbb{I}(s=s_{t+1}^{j}),\quad j,l=1,...,n;\ t=0,1,...,\infty\\ &\mathcal{R}_{i}(s_{t}^{l},a_{j})=f(t)\mathbb{I}(i=j),\quad i,j,l=1,...,n;\ t=0,1,...\infty\\ &f(t)=\left\{\begin{array}[]{ll}A,&t\leq n-1\\ B,&t\geq n\\ \end{array}\right.\end{split}$ (23)
@@
$\begin{split}J_{\mathcal{M}_{i}}^{*}=&A+A\gamma+...+A\gamma^{n-1}+B\gamma^{n}+B\gamma^{n+1}+...\\ =&A\frac{1-\gamma^{n}}{1-\gamma}+B\frac{\gamma^{n+1}}{1-\gamma}=1\\ J_{\mathcal{T}}^{*}=&\frac{1}{n}\sum_{i=1}^{n}J_{\mathcal{M}_{i}}^{*}=1.\\ \end{split}$ (24)
@@
$\begin{split}J_{\mathcal{T}}(\pi)=&\frac{1}{n}(A+A\gamma+...+A\gamma^{n-1}+B\gamma^{n}+B\gamma^{n+1}+...)\\ =&\frac{1}{n},\\ J_{\mathcal{T}}^{1}=&J_{\mathcal{T}}^{2}=\max_{\pi\in\Pi_{2}}J_{\mathcal{T}}(\pi)=\frac{1}{n}\leq\epsilon_{1}.\end{split}$ (25)
@@
$\begin{split}\hat{\pi}(a_{j}|\hat{\tau}_{t})=&\mathbb{I}(j=t+1),\quad t=0,1,...,n-1\\ \hat{\pi}(a_{j}|\hat{\tau}_{t})=&\mathbb{I}(j=i),\quad t=n,...,\infty\\ \end{split}$ (26)
@@
$\begin{split}&J_{\mathcal{T}}^{3}\geq J_{\mathcal{T}}(\hat{\pi})\\ =&\frac{1}{n}(A+A\gamma+...+A\gamma^{n-1})+B\gamma^{n}+B\gamma^{n+1}+...\\ =&\frac{A}{n}\frac{1-\gamma^{n}}{1-\gamma}+B\frac{\gamma^{n+1}}{1-\gamma}=1-\frac{(n-1)A}{n}\frac{1-\gamma^{n}}{1-\gamma}\\ \geq&1-\epsilon_{2}.\end{split}$ (27)
Thus we have proven this result. ∎
@@ A.6 Proof and Discussion of Theorem 4
In this part, we introduce an informed version of Theorem 4, about why optimizing $p(\mathcal{M}|l)\forall l\in\mathcal{L}$ is beneficial for task generalization, with detailed proofs. Recall that we consider the policy hypothesis $\mathcal{H}_{3}$ here, that each policy $\pi:\mathcal{L}\rightarrow\Delta(\mathcal{A}),\mathcal{L}=\cup_{t=1}^{\infty}\mathcal{L}_{t},\mathcal{L}_{t}=(\mathcal{S}\times\mathcal{A}\times\mathbb{R})^{t-1}\times\mathcal{S}$. As directly such $\mathcal{S}-\mathcal{A}-\mathcal{R}$ memorized policy is difficult, we consider an alternative MDP as ~$\tilde{\mathcal{M}}=(\mathcal{L},\mathcal{A},\mathcal{P}_{\mathcal{M}},\mathcal{R}_{\mathcal{M}},\gamma)$. For $\forall l=(s_{1},a_{1},r_{1},...,s_{t})\in\mathcal{L}_{t}\subseteq\mathcal{L}$, we can sample the action $a_{t}$ from the distribution $\pi(\cdot|l)$. Then the environment will feedback the reward signal $r_{t}=\mathcal{R}_{\mathcal{M}}(l,a_{t})=\mathcal{R}_{\mathcal{M}}(s_{t},a_{t})$, we can sample $s_{t+1}$ from the distribution $\mathcal{P}(\cdot|s_{t},a_{t})$, and the next environment state will be $l^{\prime}=(s_{1},a_{1},r_{1},...,s_{t},a_{t},r_{t},s_{t+1})\in\mathcal{L}_{t+1}$. In summary, we set the distribution $p(l^{\prime}|l,a_{t})$ as the corresponding dynamic $\mathcal{P}_{\mathcal{M}}$. Notice that although all $\mathcal{M}\in\mathcal{T}$ shares the same dynamic $\mathcal{P}$, their new dynamic $\mathcal{P}_{\mathcal{M}}$ are different since the dynamic is related to the given reward. An obvious advantage of introducing $\tilde{\mathcal{M}}$ is that $\pi\in\Pi_{3}$ is now Markovian in $\tilde{\mathcal{M}}$ and it is much easier to analyze its performance.
@@
$\begin{split}&J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]\\ =&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\ -\left.\int_{a,\mathcal{M}}\pi(a|l)p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)dad\mathcal{M}\right]dl\\ \geq&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right.\\ &\qquad\qquad\quad-\left.\max_{a}\int_{\mathcal{M}}p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right]dl,\end{split}$ (28)

### v5

@@ A.3 The proof of Theorem 2
[…]
$\begin{split}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=&\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{\tau\sim\mathcal{P},\pi}[R_{\mathcal{M}}(\tau)]=\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}^{\mathcal{M}}\right]=\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[r_{t}^{\mathcal{M}}]\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[\mathcal{R}_{\mathcal{M}}(s_{t}^{\mathcal{M}},a_{t}^{\mathcal{M}})]\right]=\mathbb{E}_{\tau\sim\mathcal{P},\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}[\bar{\mathcal{R}}(s_{t}^{\mathcal{M}},a_{t}^{\mathcal{M}})]\right]\\ =&\mathbb{E}_{\tau\sim\mathcal{P},\pi}[R_{\bar{\mathcal{M}}}(\tau)]=J_{\bar{\mathcal{M}}}(\pi).\end{split}$ (14)
@@
$\begin{split}J_{\mathcal{T}}^{2}&=\max_{\pi\in\Pi_{2}}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=\max_{\pi\in\Pi_{2}}J_{\bar{\mathcal{M}}}(\pi)=\max_{\pi\in\Pi_{1}}J_{\bar{\mathcal{M}}}(\pi)=\max_{\pi\in\Pi_{1}}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}[J_{\mathcal{M}}(\pi)]=J_{\mathcal{T}}^{1}\\ &\leq\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[\max_{\pi\in\Pi_{1}}J_{\mathcal{M}}(\pi)\right]=J_{\mathcal{T}}^{*}.\end{split}$ (15)
Thus we have proven this result.
@@ A.4 The Proof of Theorem 3
Proof A.3.
Our proof follows some previous work [53, 54]. First, we consider the bellman equation of the value function of $\pi,\pi_{\mathcal{M}}^{*}\in\Pi_{1}$ in $\mathcal{M}$ as
@@
$\begin{split}&V_{\mathcal{M},\pi}(s)-V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)\\ =&\gamma\sum_{a}\Delta\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s^{\prime})+\gamma\sum_{a}\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)\Delta V(s^{\prime})+\sum_{a}\Delta\pi(a|s)\mathcal{R}(s,a)\\ =&\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)+\gamma\sum_{a}\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)\Delta V(s^{\prime}),\end{split}$ (16)
@@
$\begin{split}\sum_{s}d_{\mathcal{M}}^{\pi}(s)\Delta V(s)=&\sum_{s}d_{\mathcal{M}}^{\pi}(s)[V_{\mathcal{M},\pi}(s)-V_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s)]\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)+\gamma\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\pi(a|s)\sum_{s^{\prime}}\mathcal{P}(s^{\prime}|s,a)\Delta V(s^{\prime})\\ =&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)+\sum_{s^{\prime}}\Delta V(s^{\prime})\left[\gamma\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\pi(a|s)\mathcal{P}(s^{\prime}|s,a)\right].\end{split}$ (17)
@@
$\begin{split}\sum_{s}d_{\mathcal{M}}^{\pi}(s)&\Delta V(s)=\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)+\sum_{s^{\prime}}\Delta V(s^{\prime})\left[d_{\mathcal{M}}^{\pi}(s^{\prime})-(1-\gamma)\mathcal{P}(s_{0}=s^{\prime})\right].\end{split}$ (18)
@@
$\begin{split}(1-\gamma)\sum_{s^{\prime}}\Delta V(s^{\prime})\mathcal{P}(s_{0}=s^{\prime})=&\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a),\end{split}$ (19)
@@
$\begin{split}J_{\mathcal{M}}(\pi)-J_{\mathcal{M}}(\pi_{\mathcal{M}}^{*})=&\sum_{s^{\prime}}\Delta V(s^{\prime})\mathcal{P}(s_{0}=s^{\prime})=\frac{1}{1-\gamma}\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}\Delta\pi(a|s)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ =&\frac{1}{1-\gamma}\sum_{s}d_{\mathcal{M}}^{\pi}(s)\sum_{a}[\pi(a|s)-\pi_{\mathcal{M}}^{*}(a|s)]Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ =&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M}}^{\pi}}\mathbb{E}_{a\sim\pi(\cdot|s)}\left(1-\frac{\pi_{\mathcal{M}}^{*}(a|s)}{\pi(a|s)}\right)Q_{\mathcal{M},\pi_{\mathcal{M}}^{*}}(s,a)\\ =&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M}}^{\pi}}\int_{\mathcal{A}}\pi(a|s)\left(1-\frac{\pi_{\mathcal{M}}^{*}(a|s)}{\pi(a|s)}\right)Q_{\mathcal{M}}^{*}(s,a)\mathrm{d}a\\ =&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M}}^{\pi}}\left[\int_{a}\pi(a|s)Q_{\mathcal{M}}^{*}(s,a)da-\max_{a}Q_{\mathcal{M}}^{*}(s,a)\right].\end{split}$ (20)
@@
$\begin{split}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi_{\mathcal{M}}^{*})-J_{\mathcal{M}}(\pi)\right]=&\frac{1}{1-\gamma}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}(\cdot)}\left[\max_{a}Q_{\mathcal{M}}^{*}(s,a)-\int_{a}\pi(a|s)Q_{\mathcal{M}}^{*}(s,a)da\right]\\ \geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}(\cdot)}\left[\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\max_{a}Q_{\mathcal{M}}^{*}(s,a)-\max_{a}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}Q_{\mathcal{M}}^{*}(s,a)\right]\\ =&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi}}\left[D_{\text{TDR}}(\mathcal{T},s)\right].\end{split}$ (21)
@@
$\begin{split}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\max_{\pi\in\Pi_{1}}J_{\mathcal{M}}(\pi)-\max_{\pi\in\Pi_{1}}\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}J_{\mathcal{M}}(\pi)\geq&\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{\mathcal{M},\pi^{*}}}\left[D_{\text{TDR}}(\mathcal{T},s)\right],\end{split}$ (22)
Thus we have proven this result.
@@ A.5 Expressive ability of $\Pi_{3}$
[…]
@@
[…]
@@
$\begin{split}J_{\mathcal{T}}(\pi)=&\frac{1}{n}(A+A\gamma+...+A\gamma^{n-1}+B\gamma^{n}+B\gamma^{n+1}+...)=\frac{1}{n},\\ J_{\mathcal{T}}^{1}=&J_{\mathcal{T}}^{2}=\max_{\pi\in\Pi_{2}}J_{\mathcal{T}}(\pi)=\frac{1}{n}\leq\epsilon_{1}.\end{split}$ (26)
@@
$\begin{split}\hat{\pi}(a_{j}|\hat{\tau}_{t})=&\mathbb{I}(j=t+1),\quad t=0,1,...,n-1\\ \hat{\pi}(a_{j}|\hat{\tau}_{t})=&\mathbb{I}(j=i),\quad t=n,$$\infty\\ \end{split}$ (27)
@@
$\begin{split}J_{\mathcal{T}}^{3}\geq J_{\mathcal{T}}(\hat{\pi})=&\frac{1}{n}(A+A\gamma+...+A\gamma^{n-1})+B\gamma^{n}+B\gamma^{n+1}+...\\ =&\frac{A}{n}\frac{1-\gamma^{n}}{1-\gamma}+B\frac{\gamma^{n+1}}{1-\gamma}=1-\frac{(n-1)A}{n}\frac{1-\gamma^{n}}{1-\gamma}\geq 1-\epsilon_{2}.\end{split}$ (28)
Thus we have proven this result.
@@ A.6 Proof and discussion of Theorem 4
In this part, we introduce an informed version of Theorem 4, about why optimizing $p(\mathcal{M}|l)\forall l\in\mathcal{L}$ is beneficial for task generalization, with detailed proofs. Recall that we consider the policy hypothesis $\mathcal{H}_{3}$ here, that each policy $\pi:\mathcal{L}\rightarrow\Delta(\mathcal{A}),\mathcal{L}=\cup_{t=1}^{\infty}\mathcal{L}_{t},\mathcal{L}_{t}=(\mathcal{S}\times\mathcal{A}\times\mathbb{R})^{t-1}\times\mathcal{S}$. As directly such $\mathcal{S}-\mathcal{A}$-$\mathcal{R}$ memorized policy is difficult, we consider an alternative MDP as $\tilde{\mathcal{M}}=(\mathcal{L},\mathcal{A},\mathcal{P}_{\mathcal{M}},\mathcal{R}_{\mathcal{M}},\gamma)$. For $\forall l=(s_{1},a_{1},r_{1},...,s_{t})\in\mathcal{L}_{t}\subseteq\mathcal{L}$, we can sample the action $a_{t}$ from the distribution $\pi(\cdot|l)$. Then the environment will feedback the reward signal $r_{t}=\mathcal{R}_{\mathcal{M}}(l,a_{t})=\mathcal{R}_{\mathcal{M}}(s_{t},a_{t})$, we can sample $s_{t+1}$ from the distribution $\mathcal{P}(\cdot|s_{t},a_{t})$, and the next environment state will be $l^{\prime}=(s_{1},a_{1},r_{1},...,s_{t},a_{t},r_{t},s_{t+1})\in\mathcal{L}_{t+1}$. In summary, we set the distribution $p(l^{\prime}|l,a_{t})$ as the corresponding dynamic $\mathcal{P}_{\mathcal{M}}$. Notice that although all $\mathcal{M}\in\mathcal{T}$ share the same dynamic $\mathcal{P}$, their new dynamic $\mathcal{P}_{\mathcal{M}}$ are different since the dynamic is related to the given reward. An obvious advantage of introducing $\tilde{\mathcal{M}}$ is that $\pi\in\Pi_{3}$ is now Markovian in $\tilde{\mathcal{M}}$, and it is much easier to analyze its performance.
@@
$\begin{split}J_{\mathcal{T}}^{*}-\mathbb{E}_{\mathcal{M}\sim\mathcal{T}}\left[J_{\mathcal{M}}(\pi)\right]=&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}-\int_{a,\mathcal{M}}\pi(a|l)p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)dad\mathcal{M}\right]dl\\ \geq&\frac{1}{1-\gamma}\int_{\mathcal{L}}p(l)\left[\int p(\mathcal{M}|l)\max_{a}Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}-\max_{a}\int_{\mathcal{M}}p(\mathcal{M}|l)Q_{\mathcal{M}}^{*}(l,a)d\mathcal{M}\right]dl,\end{split}$ (29)